Consider four mutually exclusive alternatives, each having an 8-year

useful life: If the minimum attractive rate of return is 8%, which

alternative should be selected (use IRR incremental)?

A B C D

First cost $1500 $100 $900 $500

Annual Benefit $122 $120 $97 $122

Salvage value $750 $500 $500



1
Expert's answer
2021-06-21T12:20:44-0400

For A:


"(1500-750)\times(1+0.08)^8+8\times122=2364"

B:


"(100-500)\times(1+0.08)^8+8\times120=220"

C:


"(900-500)\times(1+0.08)^8+8\times97=1516"

D:


"(500-0)\times(1+0.08)^8+8\times122=1901"

Answer: A